# Who Wrote this Code? Watermarking for Code Generation

### v3

@@ Appendix F Further Pareto Frontier Results on StarCoder/LLaMA2
Figure 6: The tradeoff between AUROC and pass@1 of detecting real and generated samples of HumanEval and MBPP datasets. The pink line represents a Pareto frontier of SWEET, while the blue line represents that of WLLM. In all tasks and the entropy threshold configurations, SWEET shows consistent dominance. The red/orange line and circles are the points used in Table 1.
Figure 7: Effect of general prompts in SWEET in HumanEval. In this setting, the detector does not know what information would have been included in a prompt if the given sample source code had been model-generated. SWEET appends the sample to the fixed number of ’general prompts’ that contain no information except for the format consistent with the answer. The purple line represents the Pareto frontier of the ’General prompts’ version SWEET. Our approaches with general prompts still outperform WLLM in both code quality preservation and watermark detection, drawing the Pareto frontiers ahead of those of WLLM.
@@
Appendix G More Details about Experiments with General Prompts
All general prompts we mentioned in Sec 6.2 at HumanEval task are listed below: These prompts are chosen randomly without any prompt tuning. def solution(*args):     """     Generate a solution     """ <filename>solutions/solution_1.py # Here is the correct implementation of the code exercise def solution(*args): def function(*args, **kargs):     """     Generate a code given the condition     """ from typing import List def my_solution(*args, **kargs):     """     Generate a solution     """ def foo(*args):     """     Solution that solves a problem     """
Appendix H Analysis of Computation Cost
@@
WLLM does not require any additional computation as it only needs a random number generator and a seed number to put. On the other hand, all zero-shot post-hoc detection methods excluding DetectGPT need at least one forward pass of that LLM. DetectGPT needs to run forward passes as much as the number of perturbations for increased accuracy (the original paper generated 100 perturbed samples, so we did the same). Our method needs one time forward pass to calculate the entropy, which is the same with zero-shot post-hoc detection methods except for DetectGPT. However, we demonstrated that our method outperforms baselines even when utilizing a smaller surrogate model (Sec 6.3), indicating the capability of computationally more efficient employment. On the other hand, while EXP-edit does not need LLM for detecting watermarks, it requires measuring the Levenshtein distance to compute the test statistic, which demands an extensive calculation of $O(mnk^{2})$, where $m$ be the length of the target text, $n$ be the length of the watermark key sequence, and $k$ be the block size. Moreover, $T=500$ times of test statistic is also necessary for reporting the p-value. Although these computations do not require LLM and can be implemented in parallel, one can consider the computation cost of EXP-edit as high.
Figure 10: Watermark detection performance on renamed variables in the code. For each watermark method, we choose 273 source codes from the MBPP task, for which three methods succeed in generating with no syntax error. We set $\gamma=0.25$ and $\delta=3.0$ for WLLM and SWEET. For EXP-edit, we search the hyperparameter for the block size in [20,30,40] with a high entropy setting. We use five random seeds for renaming and calculate the average AUROC scores.
Appendix I Analysis of Lexical Type Distributions
@@
I.1 List of Lexical Types
@@
I.2 Lexical Types Distributions Above Threshold
@@
I.3 Lexical Types Distributions Below Threshold
@@
Appendix J Further Analysis of Breakdown of Post-hoc methods

### v4

@@ Appendix F Further Pareto Frontier Results on StarCoder/LLaMA2
Figure 6: The tradeoff between AUROC and pass@1 of detecting real and generated samples of HumanEval, MBPP, and DS1000 datasets. The pink line represents a Pareto frontier of SWEET, while the blue line represents that of WLLM. In all tasks and the entropy threshold configurations, SWEET shows consistent dominance. The red/orange line and circles are the points used in Table 1.
Figure 7: The tradeoff between AUROC and pass@1 of detecting real and generated samples of C++ and Java of HumanEvalPack datasets. The pink line represents a Pareto frontier of SWEET, while the blue line represents that of WLLM. In all tasks and the entropy threshold configurations, SWEET shows consistent dominance. The red/orange line and circles are the points used in Table 2.
Figure 8: Effect of general prompts in SWEET in HumanEval. In this setting, the detector does not know what information would have been included in a prompt if the given sample source code had been model-generated. SWEET appends the sample to the fixed number of ‘general prompts’ that contain no information except for the format consistent with the answer. The purple line represents the Pareto frontier of the ‘General prompts’ version SWEET. Our approaches with general prompts still outperform WLLM in both code quality preservation and watermark detection, drawing the Pareto frontiers ahead of those of WLLM.
@@
Appendix G Detectability with Varying Code Lengths
Appendix H Entropy Threshold Calibration
Figure 12: The relationship between the entropy threshold and the pseudo-metric for z-score, $z^{\prime}$ in Eq. 8, calculated based on CodeSearchNet dataset. The blue region $[0.820,0.871]$ is the best entropy threshold estimated by the calibration method described in Appendix H.
This section proposes a method to calibrate the best entropy threshold effectively. To detect machine-generated texts, the z-score of them should be high. Let’s say there is a watermarked text sequence with an entropy threshold $\tau$, and the length is $N$. The z-score is calculated as in Equation 1. Thus $z\propto\sqrt{\frac{N^{h}}{N}}(\frac{N^{h}_{G}}{N^{h}}-\gamma)$, assuming a fixed $N$. If we can find a relationship between $\tau$ and $\frac{N^{h}}{N}$, and $\tau$ and $\frac{N^{h}_{G}}{N^{h}}$, we could choose $\tau$ that maximizes the z-score in a fixed $N$. We denote $z^{\prime}$ as a pseudo-metric for estimating z-score:
$z\textquoteright=E[\frac{N^{h}}{N}](E[\frac{N^{h}_{G}}{N^{h}}]-\gamma)$ (8)
With logits generated by an LLM, we can calculate entropy $H$ and the probability of sampling a green token after adding $\delta$ to green tokens’ logits. We call the expectation of it over the randomness of green/red list partitioning as $P_{G}$. We model the distribution of logits that LLM generates as a probability distribution function $P(H,P_{G};\gamma,\delta)$. We approximate the tokens in a text sequence as i.i.d, then we can write as follows:
$E[\frac{N^{h}}{N}]=P(H>\tau,P_{G})$
$E[\frac{N^{h}_{G}}{N^{h}}]=E[P_{G}|H>\tau]$
To estimate $P(H,P_{G};\gamma,\delta)$ in the Python language domain of StarCoder, we use a code corpus, CodeSearchNet (Husain et al., 2019). Specifically, we feed the Python corpus of CodeSearchNet to our model and obtain all logits for each time step and calculate $H$ and $P_{G}$. For $P_{G}$, we averaged the probability of sampling a green token from a watermarked distribution with 500 random green/red list partitions. We regard the pair ($H$, $P_{G}$) per one logits as an unnormalized joint discrete distribution.
The results are presented in Figure 12. The $z\textquoteright$ ’ is the highest when the entropy threshold is in $[0.820,0.871]$. It aligns with the results in Figure 4, where the optimal threshold value lies around 0.3$\sim$0.9. Therefore, the threshold value found here is a good starting point for searching for the optimal threshold value. The computational cost is only one forward pass across the corpus we used. The result indicates that we can use the information in the code corpus to calibrate an entropy threshold effectively.
Appendix I Analysis of Computation Cost
@@
WLLM does not require any additional computation as it only needs a random number generator and a seed number to put. On the other hand, all zero-shot post-hoc detection methods excluding DetectGPT need at least one forward pass of that LLM. DetectGPT needs to run forward passes as much as the number of perturbations for increased accuracy (the original paper generated 100 perturbed samples, so we did the same). Our method needs one time forward pass to calculate the entropy, which is the same with zero-shot post-hoc detection methods except for DetectGPT. However, we demonstrated that our method outperforms baselines even when utilizing a smaller surrogate model (Sec 6.3), indicating the capability of computationally more efficient employment. On the other hand, while EXP-edit does not need LLM for detecting watermarks, it requires measuring the Levenshtein distance to compute the test statistic. Specifically, it demands an extensive calculation of $O(mnk^{2})$, where $m$ be the length of the target text, $n$ be the length of the watermark key sequence, and $k$ be the block size. Moreover, $T=500$ times of test statistic is also necessary for reporting the p-value. Although these computations do not require LLM and can be implemented in parallel, one can consider the computation cost of EXP-edit as high.
Appendix J Analysis of Lexical Type Distributions
@@
J.1 List of Lexical Types
@@
J.2 Lexical Types Distributions Above Threshold
@@
J.3 Lexical Types Distributions Below Threshold
@@
Appendix K Further Analysis of Breakdown of Post-hoc methods